# Front stability of infinitely steep travelling waves in population biology

### 21 Dec 2023 v1

@@ 4 Conclusion and future work
with $\delta c=1\times 10^{-6}$. This method for computing the solution of the $\mathcal{O}(1)$ boundary value problem requires that we specify $\kappa$ and then estimate $c$ iteratively by shooting. The efficiency of this approach strongly depends upon providing a good initial estimate of $c$ for our choice of $\kappa$ and $m$, and there are many ways to make a good estimate. One way is to use the phase plane tools described in Section 3.2 to calculate an accurate estimate of $c$ for a particular choice of $m$ and $\kappa$. A far simpler option is to use Equation (14) to estimate $c$ for a given value of $\kappa$. Note that Equation (14) is exact for $m=1$ and $|c|\ll 1$ only, however in our experience this relationship still provides a reasonable initial estimate of $c$ that is improved via shooting.
@@
To solve for the $\mathcal{O}(\varepsilon)$ term in the linear stability analysis we re–write Equations (20a)–(20d) in terms of $\theta_{1}=u_{1}^{(1+m)}$, and then consider the resulting boundary value problem on an appropriately truncated domain, which gives,
$\left(\frac{\mathrm{d}^{2}\theta_{1}}{\mathrm{d}\xi^{2}}\right)u_{0}^{m+1}-c\theta_{1}\;m\left(\frac{\mathrm{d}u_{0}}{\mathrm{d}\xi}\right)-\omega\theta_{1}u_{0}+cu_{0}\left(\frac{\mathrm{d}\theta_{1}}{\mathrm{d}\xi}\right)-$
$u_{0}\left(\theta_{1}\left(\left(q^{2}u_{0}^{m-1}+2\right)u_{0}-1\right)\right)=-u_{0}^{m+1}\left(\frac{\mathrm{d}u_{0}}{\mathrm{d}\xi}\right)\left(\omega+q^{2}u_{0}^{m}\right)\;\;\text{on}\;\;-\xi_{\textrm{max}}<\xi<0,$ (30a)
$\theta_{1}=0\;\;\text{at}\;\;\xi=-\xi_{\textrm{max}},$ (30b)
$\theta_{1}=0,\;\;\text{at}\;\;\xi=0,$ (30c)
$\frac{\mathrm{d}\theta_{1}}{\mathrm{d}\xi}$ $=-\frac{\omega}{\kappa}\;\;\text{at}\;\;\xi=0.$ (30d)
@@
$\theta_{1}^{(0)}=0,$ (31a)
$\begin{gathered}(u_{0}^{(i)})^{(m+1)}\left(A\;\theta_{1}^{(i+1)}+B\;\theta_{1}^{(i)}+C\;\theta_{1}^{(i-1)}\right)+c\;u_{0}^{(i)}\;\frac{\theta_{1}^{(i+1)}-\theta_{1}^{(i-1)}}{h_{i}^{+}+h_{i}^{-}}-c\;\theta_{1}^{(i)}\;\frac{u_{0}^{(i+1)}-u_{0}^{(i-1)}}{h_{i}^{+}+h_{i}^{-}}\\ -\omega\;\theta_{1}^{(i)}\;u_{0}^{(i)}-u_{0}^{(i)}\left(\theta_{1}^{(i)}\left(\left({(u_{0}^{(i)})}^{(m-1)}q^{2}+2\right)u_{0}^{(i)}-1\right)\right)=-(u_{0}^{(i)})^{(m+1)}\frac{u_{0}^{(i+1)}-u_{0}^{(i-1)}}{h_{i}^{+}+h_{i}^{-}}\;\left(\omega+q^{2}{u^{(i)}_{0}}^{m}\right),\end{gathered}$ (31d)
$\theta_{1}^{(N)}=0,$ (31e)
where the superscript $i=2,\dots,N-1$ refers to the $i$th internal mesh point, and the coefficients $A={2}/({h^{+}_{i}(h_{i}^{+}+h_{i}^{-})})$, $B={-2}/({h_{i}^{+}h_{i}^{-}})$. and $C={2}/({h^{-}_{i}(h_{i}^{+}+h_{i}^{-})})$ arise by applying standard finite difference stencils on the nonuniform mesh McCue2022 . Values of $c$ and $u_{0}$ are known inputs that we have previously calculated as a result of solving the $\mathcal{O}(1)$ boundary value problem. Similarly, the wavenumber $q$ is also a specified input. Here we can calculate $\theta_{1}$ by noting that Equations (31a)–(31e) are linear in $\theta_{1}^{(i)}$ so we solve this system using Julia’s backslash operator.
@@
$g^{(k)}=\omega^{(k)}+\kappa\left(P\theta_{1}^{({N-2},k)}-Q\theta_{1}^{({N-1},k)}+R\theta_{1}^{({N},k)}\right),$ (32)
where $P=h^{+}_{N-1}/({h^{-}_{N-1}(h^{+}_{N-1}+h^{-}_{N-1})})$, $Q=(h^{+}_{N-1}+h^{-}_{N-1})/(h^{+}_{N-1}h^{-}_{N-1})$,, and $R=(2h^{+}_{N-1}+h^{-}_{N-1})/({h^{+}_{N-1}(h^{+}_{N-1}+h^{-}_{N-1})})$. To satisfy Equation (30d) we update our estimate of the growth rate so that $g=0$ using Newton-Raphson iteration
@@
where we estimate $g^{\prime}$ using a finite difference approximation

### 06 Jan 2024 v2

@@ 4 Conclusion and future work
with $\delta c=1\times 10^{-6}$. This method for computing the solution of the $\mathcal{O}(1)$ boundary value problem requires that we specify $\kappa$ and then estimate $c$ iteratively by shooting. The efficiency of this approach strongly depends upon providing a good initial estimate of $c$ for our choice of $\kappa$ and $m$, and there are many ways to make a good estimate. One way is to use the phase plane tools described in Section 3.2 to calculate an accurate estimate of $c$ for a particular choice of $m$ and $\kappa$. A much simpler option is to use Equation (14) to estimate $c$ for a given value of $\kappa$. Note that Equation (14) is exact for $m=1$ and $|c|\ll 1$ only, however this relationship still provides a reasonable initial estimate of $c$ that is then iteratively improved by shooting.
@@
To solve for the $\mathcal{O}(\varepsilon)$ term in the linear stability analysis we re–write Equations (20a)–(20d) in terms of $\phi_{1}=u_{0}^{m}u_{1}$, and then consider the resulting boundary value problem on an appropriately truncated domain, which gives,
$\left(\frac{\mathrm{d}^{2}\phi_{1}}{\mathrm{d}\xi^{2}}\right)u_{0}^{m+1}-c\phi_{1}\;m\left(\frac{\mathrm{d}u_{0}}{\mathrm{d}\xi}\right)-\omega\phi_{1}u_{0}+cu_{0}\left(\frac{\mathrm{d}\phi_{1}}{\mathrm{d}\xi}\right)$
$-\phi_{1}\left(q^{2}u_{0}^{m+1}+2u_{0}^{2}-u_{0}\right)=-u_{0}^{m+1}\left(\frac{\mathrm{d}u_{0}}{\mathrm{d}\xi}\right)\left(\omega+q^{2}u_{0}^{m}\right)\;\;\text{on}\;\;-\xi_{\textrm{max}}<\xi<0,$ (30a)
$\phi_{1}=0\;\;\text{at}\;\;\xi=-\xi_{\textrm{max}},$ (30b)
$\phi_{1}=0\;\;\text{at}\;\;\xi=0,$ (30c)
$\frac{\mathrm{d}\phi_{1}}{\mathrm{d}\xi}$ $=-\frac{\omega}{\kappa}\;\;\text{at}\;\;\xi=0.$ (30d)
@@
$\phi_{1}^{(0)}=0,$ (31a)
$\begin{gathered}\left(u_{0}^{(i)}\right)^{(m+1)}\left(A\phi_{1}^{(i+1)}+B\phi_{1}^{(i)}+C\phi_{1}^{(i-1)}\right)+cu_{0}^{(i)}\dfrac{\phi_{1}^{(i+1)}-\phi_{1}^{(i-1)}}{h_{i}^{+}+h_{i}^{-}}-c\theta_{1}^{(i)}\dfrac{u_{0}^{(i+1)}-u_{0}^{(i-1)}}{h_{i}^{+}+h_{i}^{-}}\\ -\omega\;\phi_{1}^{(i)}\;u_{0}^{(i)}-\phi_{1}^{(i)}\left({(u_{0}^{(i)})}^{(m+1)}q^{2}+2{(u_{0}^{(i)})}^{2}-u_{0}^{(i)}\right)=-\left(u_{0}^{(i)}\right)^{(m+1)}\dfrac{u_{0}^{(i+1)}-u_{0}^{(i-1)}}{h_{i}^{+}+h_{i}^{-}}\left(\omega+q^{2}{u^{(i)}_{0}}^{m}\right),\end{gathered}$ (31d)
$\phi_{1}^{(N)}=0,$ (31e)
where the superscript $i=2,\dots,N-1$ refers to the $i$th internal mesh point, and the coefficients $A={2}/({h^{+}_{i}(h_{i}^{+}+h_{i}^{-})})$, $B={-2}/({h_{i}^{+}h_{i}^{-}})$. and $C={2}/({h^{-}_{i}(h_{i}^{+}+h_{i}^{-})})$ arise by applying standard finite difference stencils on the nonuniform mesh McCue2022 . Values of $c$ and $u_{0}$ are known inputs that we have previously calculated as a result of solving the $\mathcal{O}(1)$ boundary value problem. Similarly, the wavenumber $q$ is also a specified input. Here we note that Equations (31a)–(31e) are linear in $\phi_{1}^{(i)}$ so we solve this linear system using Julia’s backslash operator.
@@
$g^{(k)}=\omega^{(k)}+\kappa\left(P\phi_{1}^{({N-2},k)}-Q\phi_{1}^{({N-1},k)}+R\phi_{1}^{({N},k)}\right),$ (32)
where $P=h^{+}_{N-1}/({h^{-}_{N-1}(h^{+}_{N-1}+h^{-}_{N-1})})$, $Q=(h^{+}_{N-1}+h^{-}_{N-1})/(h^{+}_{N-1}h^{-}_{N-1})$, and $R=(2h^{+}_{N-1}+h^{-}_{N-1})/({h^{+}_{N-1}(h^{+}_{N-1}+h^{-}_{N-1})})$. To satisfy Equation (30d) we update our estimate of the growth rate so that $g=0$ using Newton-Raphson iteration
@@
$\lVert\omega^{(k+1)}-\omega^{(k)}\rVert<1\times 10^{-6}$. To implement this algorithm we estimate $g^{\prime}$ using a finite difference approximation,